OpenStudy (anonymous):

It is true that 49.2 kilometers are equivalent to 30.57 miles. From this information, derive an equation converting distance in km to distance in miles. What is the distance in miles of 96.2 km?

OpenStudy (anonymous):

given: 49.2ki =30.57 mi ?

OpenStudy (anonymous):

\[\frac{49.2Ki}{96.2Ki}=\frac{30.57mi}{m?}\]cross and time 49.2m=46.2*30.57

OpenStudy (anonymous):

so 46.2x30.57=49.2m?

OpenStudy (anonymous):

i am so lost

OpenStudy (anonymous):

I was wrong second line

OpenStudy (anonymous):

49.2m=96.2*30.57

OpenStudy (anonymous):

m=(2.940.834)/49.2 59.77304878=miles than round

OpenStudy (anonymous):

so 60 miles?

OpenStudy (anonymous):

Since you have that 49.2 km=30.57 mi Since you need to find distances in miles, you divide by 30.57 both sides and you will have: 1.609km=1mi or 1.609km/mi That's part 1 of your question Now, to find the distance in miles, just divide the distance in kilometers, by 1.609 and you will get the distance in miles.

OpenStudy (anonymous):

yes, round